OpenStudy (anonymous):

Use quotient rule and simplify. Write your answer in positive exponent. 1.) 8^w 8^6/ 8^w 8^2 / meaning over or divide Please show work and thank you.

OpenStudy (p0sitr0n):

8^w8^6 / 8^w8^2=8^6 / 8^2 = 8^4

OpenStudy (anonymous):

Thank you. But may you please show me how you got that I have to show work.